# VSF: Simple, Efficient, and Effective Negative Guidance in Few-Step Image Generation Models By Value Sign Flip

### v2

@@ 3.2 Attention Masking and Duplication of Negative Embedding
The above method works well for cross-attention-based methods, where attention only exists between image-to-image in self-attention layers and image-to-text in cross-attention layers. However, it requires modification, including masking and duplication, to work in MMDiT-style models such as SD3.5 (sd, 3), where all image and text tokens are concatenated into a single sequence before attention.
In the standard MMDiT-style setup without our guidance, the sequence inputs for the attention module are:
$\mathbf{I},\ \mathbf{P}]\ \text{and}\ [\mathbf{I},\ \mathbf{N}]$
If we concatenate all tokens into a single sequence without any modification, we will get:
$\mathbf{I},\ \mathbf{P},\ \mathbf{N}]$
where $\mathbf{I}$ represents image tokens, $\mathbf{P}$ represents positive prompt tokens, and $\mathbf{N}$ is the negative prompt. During attention, queries, keys, and values are all projected from this combined sequence.
If we apply a sign flip to the negative prompt values by scaling $V_{\mathbf{N}}$ with $\alpha$ (where $V$ is the value projection, this flipped content affects all attention paths involving $V_{\mathbf{N}}$. That includes not only the intended interaction between image and negative prompt $(\mathbf{I}\rightarrow\mathbf{N})$, but also undesired interactions such as positive-to-negative $(\mathbf{P}\rightarrow\mathbf{N})$ and negative-to-negative $(\mathbf{N}\rightarrow\mathbf{N})$ (which the value will cancel itself). These unintended interactions can distort the behavior of the model since the flipped signal influences more than just the image.
To address this, we introduce a duplication of the negative prompt. One copy remains unflipped and unscaled, denoted $\mathbf{N}^{(0)}$, and the value (and only value) of other is flipped and scaled, denoted $V_{\mathbf{N}^{(1)}}=-\alpha\cdot V_{\mathbf{N}}$. The sequence becomes:
$\mathbf{I},\ \mathbf{P},\ \mathbf{N}^{(0)},\ \mathbf{N}^{(1)}]$
@@
Queries are sourced from $\mathbf{I}$, $\mathbf{P}$, and $\mathbf{N}^{(0)}$, while keys and values include all four components.
With a similar idea in Wang et al. (2025), we apply attention masks to isolate the effect of the flipped negative prompt. Specifically, $\mathbf{N}^{(0)}$ is only allowed to attend to $\mathbf{I}$ and to itself, while $\mathbf{N}^{(1)}$ is only attended to by $\mathbf{I}$. Since $\mathbf{N}^{(1)}$ does not act as a key or value in any attention query, it produces no associated output; instead, $\mathbf{N}^{(0)}$ serves as the effective negative prompt tokens passed to the subsequent MLP layer and into the next attention layer, where it will be flipped again. Note that $N^{(0)}$ used as queries, keys, and values while $N^{(1)}$ is only used as keys and values.
This setup allows updates to the negative prompt based on attention from the image and from itself, while keeping the unflipped form active in the MLP path. It also prevents interference between positive and negative prompts and ensures that the flipped negative content affects only the intended image-to-negative attention path.
3.3 Attention Bias
3.4 Padding Removel
In most models from Huggingface Diffusers (von Platen et al., ), padding tokens in the text input are typically not masked during attention. This is likely because the models have learned to ignore padding, and masking them would add unnecessary overhead (due to some attention implementations like FlashAttention-2 (Dao, 2023) that do not support arbitrary masking). However, when we invert the sign of the padding tokens, it degrades output quality. This could be because, although these tokens carry no semantic meaning, the sign-flipping introduces unseen states into the attention mechanism. To mitigate this, we remove padding tokens from the negative prompt embeddings. For the positive prompt, we retain padding tokens, as they do not introduce novel tokens and can improve generation quality. This aligns with training conditions and may allow the model to use padding positions as registers for auxiliary information.
@@ 4.1 Dataset
Following Park et al. (2025), we use ChatGPT o3 (Open AI, ) to generate pairs of prompts and negative prompts. Unlike prior work, our prompts are intentionally more challenging: the negative prompt is typically related to the positive one, and as a critical component—e.g., the positive prompt of a bike could have a negative prompt of “wheels”. Additional examples are shown in Figure 7. Besides prompts, two questions are generated at the same time for later evaluation, one query if the image has the main object, either with or without the negative element, and the other one queries if the negative prompt element is missing. Prompts are generated in batches. Due to the fact that the model might output similar concepts for different prompts, it may introduce some repetition across batches or within batches with different phasing. There are 200 prompts generated, and we run them with 2 different seeds for the main results.
4.2 Baseline
@@
We chose NAG (Chen et al., 2025a) and NASA Nguyen et al. (2024) as our baseline. We also used a base model without negative guidance as a bare baseline, aiming to show the lower bound of the dataset (i.e., how likely the positive prompt will introduce the negative concept, if there is no negative guidance). Because NASA’s original source code was not publicly available at the time of writing, we reimplemented it based on NAG’s codebase. Specifically, we replaced the guidance equation from NAG (Eq. 4) with NASA’s equation (Eq. 3), removed normalization and blending, and enabled guidance when the scale is greater than 0 (instead of 1). Additionally, to compare our method in few-step models with CFG on original non-few-step models, we also used the original Stable Diffusion 3.5 Large with CFG as a baseline.
4.3 Metric
Following Park et al. (2025); Wei et al. (2025), we used multimodal large language models (MLLM), specifically llama-4-maverick-17b-128e-instruct-fp8 (llama), to evaluate if the generated image follows the positive prompt and the negative prompt using the two questions generated during prompt generation. LLaMA 4 Maverick has a very high image reasoning MMMU mmmu score, higher than Gemma 3 and even GPT-4o gpt4o. We avoided using the same model (o3) for both evaluation and generation for cost control and to avoid bias within a model. We did not evaluate the quality of the generated images using models like ImageReward (Xu et al., 2023) or HPSv2 Wu et al. (2023) as in NASA or NAG, as current quality or human preference assessment models do not account for negative prompts. Removing a key element from the positive prompt (e.g., removing the roof from a house) is likely to reduce perceived quality, since the result deviates from what is considered “normal,” even though that is the intended outcome. Both ImageReward and HPSv2 are built on top of image-text alignment models (CLIP (Radford et al., 2021) or BLIP (Li et al., 2022)), which will likely lead to a decreased score when the main object is missing a critical part. Thus, we also let the MLLM rate the image quality from 0-1 for each image and told it to ignore the abnormality of following the negative prompt. Given that it is very hard to quantify the performances of methods in these cases, we used qualitative methods as the main evaluation and comparison, following Wang et al. (2025). We did observe that MLLM can make mistakes, especially when there is ambiguity or when the unwanted element is hard to see. However, we believe that in general, under 400 images, the mistakes are minor. To compensate for this, we also did a human evaluation between our model and NAG for pairwise comparsion.
4.4 Hyper-parameter Tunning
Although NAG (Chen et al., 2025a) also targeted negative concept avoidance, its primary focus was on its effects on improving generation quality (using words like “blurry” or “low quality” as a negative prompt). We believe the hyperparameters reported in their work were tuned with an emphasis on quality rather than negation handling. Therefore, we re-tuned their hyperparameters moderately and manually targeting guidance scale ($\phi$), blending factor ($\alpha$), and normalization factor ($\tau$). We will report experimental results on both original NAG (noted as NAG) and the improved hyperparameter version (noted as NAG++). The final hyperparameters used are $\phi=11,\alpha=0.5,\tau=5$. This pushes the NAG to the edge of acceptable visual quality. Although we cannot ensure this is the best performing hyperparameters, we believe this version of NAG has better negative prompt following, with sacrifices in quality. These parameters are not swept to avoid over-fitting. However, we did conduct experiments on the positive-negative-quality trade-off study later in this paper.
4.5 Results
Quantitative results from MLLM as judge evaluation are shown in Table 2, and qualitative results are shown in Figure 7. P-value for the negative score is tested with the McNemar test. Qualitative results are from 10 randomly selected samples.
It is important to highlight that the MLLM assigns relatively generous quality scores; significantly distorted images may still receive high ratings. Empirically, we observed that images scoring around 60 are heavily distorted and exhibit numerous artifacts. For example, the image on the left of Figure 6 has a quality score of 70 yet displays severe degradation, and the image in the middle has noticeable distortion but is rated with 90, and the image on the right has minor artifacts (meaning it is not perfect) while receiving a score of 100.
Figure 7: Qualitative comparison of an NAG version whose hyperparameter was tuned for negative guidance (NAG++), original NAG (NAG), and our method (VSF). These samples are randomly selected from the results.
ChatGPT said:Based on the quantitative results, VSF shows a significantly higher negative score than other methods, while maintaining comparable or better quality scores. However, its positive score is considerably lower. The trade-off between positive and negative scores is discussed in the extended section. Our method achieves a higher negative score than traditional CFG in non-few-step models, demonstrating a stronger ability to avoid negative elements even relative to the established strong baseline.
From the qualitative results, we can see that in many cases, our VSF avoided the negative prompt better, such as the plane wings and bristles examples. We can also see that in many cases, the guidance scale of NAG++ and our version of NASA is pushed to the limit of acceptable quality, yet still fails to follow the negative prompt. For VSF, the quality is between NAG and NAG++, sometimes resulting in very simple, lacking detail images like the red car, a jet without wings, or the laptop with keyboard. However, in many cases it still preserved high quality, such as the color palette, the chess, and the train. All three methods resulted in some unnatural physical formation, such as the weird location of the laptop hinge. There are also cases where all methods failed, such as removing the head from the statue and removing the nose from the teddy bear, indicating the significant challenges of our dataset. Additionally, our model sometimes removes more features than we want, such as the keyboard in the laptop and the windshield in the car, due to attention map dispersion or semantic similarity between these elements. Sometimes, VSF performs a bit worse than other methods, like in the leaves example.
4.6 Trade Off Curve
To systematically evaluate how effectively each model balanced positive prompt adherence, negative prompt adherence, and image quality, we conducted a hyperparameter sweep across each model. Specifically, we performed 66 runs for VSF and 287 runs for NAG, and 10 runs for NASA, with respect to their hyperparameter counts (2 for VSF, 3 for NAG, and 1 for NASA). A random sweep was executed besides for NASA, on which a grid search is used, and evaluations were conducted using Llama 4 (llama4) following the same criteria as previously described. Due to the large volume of runs, we limited our evaluation to the first 100 prompts with a single generation seed, potentially resulting in minor differences from earlier outcomes.
For the trade-off plot, runs were sorted by negative prompt adherence scores from highest to lowest. As we trace from the highest negative score to the lowest, we sequentially record the highest positive prompt or image quality scores encountered. Each run was marked as a critical point if it improved upon previously recorded positive prompt or quality scores. All critical points are plotted and connected, and shown in Figure 8.
Figure 8: Trade off plot of positive-negative score and quality-negative score. Both axies follows “higher is better.”
From both plots, we observe that as the negative score increases, NAG and NASA both exhibit a significantly steeper decline compared to VSF in both positive and quality scores. NASA has a steeper decline in quality compared to NAG, which is expected as NASA does not have the normalization and blending. In terms of positive score, VSF maintains scores above 90 even when the negative score rises to approximately 60. Regarding image quality, VSF similarly retains scores above 90 until a negative score of around 60, after which quality declines. In contrast, NAG and NASA both experience a sharp decline, with their quality score rapidly dropping to nearly 60 even before the negative score reaches 50.
Additionally, VSF demonstrates a broader operational range in negative scores. When necessary, it can achieve negative scores exceeding 70 while still preserving acceptable positive prompt adherence and image quality. Conversely, NAG and NASA become unacceptable in quality at negative scores below 50, limiting their practical effectiveness. Keep in mind that the MLLM usually overestimates the quality, and if an image is rated 60, it is usually completely distorted. See Figure 6 for example.
4.7 External Baselines
In addition to other guidance methods applied to SD-3.5-turbo, we evaluated several external baselines. The first baseline employs a generate-then-edit approach, loosely inspired by Generate-Plan-Edit (GraPE) (Goswami et al., 2025), but omitting the planning stage as our goal is straightforward (removing unwanted elements). Specifically, we first generated images using SD-3.5-Large-Turbo without a negative prompt, and subsequently edited out the unwanted elements using Flux Knoest (Labs et al., 2025) using prompt Remove [negative prompt].
The second baseline utilizes GPT-4o’s native image generation capability. GPT-4o has demonstrated strong prompt-following performance(Wei et al., 2025), including in handling negation tasks. As GPT-4o lacks explicit negative prompt functionality, we formatted prompts as [Positive prompt], but with no [negative prompt]. Since our focus is on evaluating negation rather than image quality, we adopted the ”low” generation setting.
The third baseline we included is Janus-4o (Chen et al., 2025b), a model distilled from GPT-4o onto the Janus-Pro base architecture (Chen et al., 2025c). Given GPT-4o’s strong prompt-following performance, we anticipated competitive results from Janus-4o. We provided negative prompts directly as negations within the positive prompts.
All measure time is measured on Colab A100 GPU and for Qwen-Image and Generate+Edit, model CPU off loading is enabled.
@@ 4.7 External Baselines
Table 1: External Baselines Comparsion
The results are presented in Table 1. The table indicates that VSF achieves the second-best negative score, only behind GPT-4o, while also demonstrating a significantly faster runtime compared to all other methods, outperforming even the generate-then-edit pipeline. The GPT-4o distilled model, Janus-4o, has an unexpectedly low negative score, which could be because they did not have enough negation-included prompts in the distillation data.
Upon visual inspection, we observed that GPT-4o effectively avoids many negative prompts, though occasionally the ambiguity within negative prompts (e.g., the term “door” referring either to the door panel or the entrance) and the vision ability of MLLM itself (i.e., hallucination) leads to lower negative scores. It is possible that the negative scores for other methods might also be underestimated. We acknowledge this as a limitation associated with using an MLLM as the evaluator.
@@ 4.8 Attention Maps
Since our proposed method performs adaptive steering based on a negative attention map, we visualize the attention maps generated during the diffusion process in Figure 4. Extracting the full attention maps is difficult because efficient implementations, such as FlashAttention, do not explicitly store these maps, and storing and computing them will require a large amount of memory. Therefore, we computed only the unnormalized attention values between the image tokens and negative prompt tokens.
Figure 4 demonstrates that when the scale is set to 0, umbrellas appear, whereas setting the scale to 3 effectively removes them. As indicated in the attention maps, image tokens corresponding to regions where umbrellas might exist (e.g., above human heads) exhibit higher attention toward the negative prompt tokens. Specifically, in steps 4 and 5, regions above the individuals on the left and right show strong negative attention, aligning with areas visually identified as umbrellas. In the final image, these highlighted regions no longer contain umbrellas, confirming that our method effectively suppresses the presence of undesired objects at specific locations.

### v5

@@ 3.2 Attention Masking and Duplication of Negative Embedding
The above method works well for cross-attention-based methods, where attention only exists between image-to-image in self-attention layers and image-to-text in cross-attention layers. However, it requires modification, including masking and duplication, to work in MMDiT-style models such as SD3.5 (Stability AI, 2024), where all image and text tokens are concatenated into a single sequence before attention.
In the standard MMDiT-style setup without our method (e.g., using CFG, NASA, or NAG), the sequence inputs for the attention module are: $\mathbf{I},\ \mathbf{P}]\ \text{and}\ [\mathbf{I},\ \mathbf{N}].$ If we concatenate all tokens into a single sequence without any modification, we will get: $\mathbf{I},\ \mathbf{P},\ \mathbf{N}],$ where $\mathbf{I}$ represents image tokens, $\mathbf{P}$ represents positive prompt tokens, and $\mathbf{N}$ is the negative prompt. During attention, queries, keys, and values are all projected from this combined sequence.
If we apply a sign flip to the negative prompt values by scaling $V_{N}=V\mathbf{N}$ with $\alpha$ (where $V$ is the value projection), this flipped content affects all attention paths involving $V_{\mathbf{N}}$. That includes not only the intended interaction between image and negative prompt $(\mathbf{I}\rightarrow\mathbf{N})$ [^1] The arrow direction is the attention direction, or the opposite direction of the information flow, but also undesired interactions such as positive-to-negative $(\mathbf{P}\rightarrow\mathbf{N})$ and negative-to-negative $(\mathbf{N}\rightarrow\mathbf{N})$ (in which the value will cancel itself). These unintended interactions can distort the behavior of the model since the flipped signal influences more than just the image.
To address this, we introduce a duplication of the negative prompt. One copy remains unflipped and unscaled, denoted by $\mathbf{N}^{(0)}$, and the value (and only value) of the other is flipped and scaled, denoted by $V_{\mathbf{N}^{(1)}}=-\alpha\cdot V_{\mathbf{N}^{(1)}}$. The sequence becomes: $\mathbf{I},\ \mathbf{P},\ \mathbf{N}^{(0)},\ \mathbf{N}^{(1)}],$ where $\mathbf{N}^{(1)}$ does not act as query in attention calculation.
@@
Additionally, inspired by Wang et al. (2025), where blocking some attention directions could improve quality, we apply attention masks to isolate the effect of $\textbf{N}^{(1)}$ to only I. Specifically, $\mathbf{N}^{(0)}$ is only allowed to attend to $\mathbf{I}$ and to itself, while $\mathbf{N}^{(1)}$ is only attended to by $\mathbf{I}$. Figure 3 shows the attention mask. Since $\mathbf{N}^{(1)}$ does not act as a key or value in any attention query, it doesn’t produce associated output. Instead, $\mathbf{N}^{(0)}$ serves as the negative prompt tokens passed to the subsequent MLP layer and into the next attention layer, where it will be flipped again. It acts as an information collector from images to collect unwanted elements and also keeps updating itself from attention to itself, matching the prompt updating in a positive prompt.
This setup allows updates to the negative prompt based on attention from the image and from itself, and keeps the unflipped form active in the MLP path. It also prevents interference between positive and negative prompts and ensures that the flipped negative content affects only the intended image-to-negative attention path.
To preserve the high quality of generated images, we also applied attention bias ($\beta$) to $\mathbf{I}\rightarrow\mathbf{N}^{(1)}$ (also shown in Figure 3) and we removed the padding tokens from the negative prompt. Details and pseudo-code of our method are in the Appendix.
@@ 4.1 Dataset
Following Park et al. (2025), we use ChatGPT o3 (Open AI, ) to generate pairs of prompts and negative prompts to construct our dataset NegGenBench. Unlike prior work, our prompts are intentionally more challenging: the negative prompt is typically related to the positive one, and as a critical component—e.g., the positive prompt of a bike could have a negative prompt of “wheels”. However, the positive prompt sometimes also uses a non-negation method to imply the item is missing, such as using terms like “empty” and “exposed” to make it more natural. Besides prompts, two questions are generated at the same time for later evaluation, one question asking if the image has the main object, either with or without the negative element, and the other one queries if the negative prompt element is missing. Prompts are generated in batches. There are 200 prompts generated, and we run them with 2 different seeds for the main results.
4.2 Baseline and Metrics
@@
We chose NAG (Chen et al., 2025a) and NASA Nguyen et al. (2024) as our baseline for few-step models. We also used a base model without negative guidance as a vanilla baseline, aiming to show the lower bound of the dataset. (i.e., how likely the positive prompt alone will help avoid negative concepts, if there is no negative guidance. This could happen either because the model is following the implication in the positive prompt, such as the word ‘missing’, or simply by chance.) Because NASA’s original source code was not publicly available at the time of writing, we reimplemented it based on NAG’s codebase. Specifically, we replaced the guidance equation from NAG (Eq. 4) with NASA’s equation (Eq. 3), removed normalization and blending, and enabled guidance when the scale is greater than 0 (instead of 1). Additionally, we compared our method in non-few-step models with CFG and used other models as external baselines (External baseline results are in Table 1, but the experiment details are in the Appendix). Since NAG was focused on quality instead of negative prompt avoidance, we re-tuned its hyperparameter such that it has stronger negation following in trade-off of quality and positive prompt following. We name this variance as NAG Strong. Same for our VSF method, we provided two different variations with different hyperparameters, focusing on quality (VSF Quality) and negative prompt following (VSF Strong). Hyperparameter details are in the Appendix.
Following Park et al. (2025); Wei et al. (2025), we used multimodal large language models (MLLM), specifically llama-4-maverick-17b-128e-instruct-fp8, to evaluate if the generated image follows the positive prompt and the negative prompt using the two questions generated during prompt generation. We did not use previous negation-aware CLIP-based work because they do not focus on missing an essential component, but simple meaning (e.g., a dog that is not on the grass). We did not use HPSv2 (Wu et al., 2023) or ImageReward (Xu et al., 2023) because they might give a low quality score for unusual objects (essential part being removed). Instead, MLLM is used to rate the image quality at the same time. At the end of our experiment, we also fine-tuned a Qwen-2.5-VL (Bai et al., 2025) model using data we generated by VSF, NAG, and NASA for better negation understanding. More details about the metrics and comparison with human validation are in the Appendix.
5 Results
Quantitative results from using LLaMA as a judge evaluation are shown in Table 2, qualitative results are shown in Figure 12 in the Appendix. Human validation is shown in Table 3. Automatic evaluation using the better negation-aware MLLM Qwen-2.5-VL is shown in Table 8 in the Appendix. Both the human validation and Qwen-2.5-VL results are aligned with our LLaMA evaluation relative ranking. It is important to highlight that the LLaMA assigns relatively generous quality scores; a score lower than 90 usually means the image already has degraded quality. Examples are in the Appendix Figure 11.
Figure 5: (Left) Style Avoidance Tests, (Right) Abstract art generated by mentioning the main object “car” in a negative prompt.
Table 1: External Baselines Comparsion
@@ 5 Results
Closed-weight Models
GPT-4o 0.978 0.705 0.954 47s
Nano Banana 0.985 0.498 0.980 14s
Based on the quantitative results, VSF Strong shows a significantly higher negative score than other methods, while maintaining comparable or better quality scores. Our more conservative method, VSF Quality, still achieved the second-highest negative score, with the highest quality score. Both VSF Strong and VSF Quality even achieve a higher negative score than traditional CFG in non-few-step models, demonstrating a stronger ability to avoid negative elements, even relative to the established strong baseline. When compared with the external baseline, VSF also gets the highest performance in open source methods and only lags behind GPT-4o and achieves comparable performance with Nano Banana.
We also tested style avoidance, as shown in Figure 5 left. Figure 5 also illustrates how our method can produce abstract art, which is typically discouraged during a model’s finetuning since reward models favor realism. This is achieved through using the same word as the main object for both positive and negative scores in VSF, as detailed in the Appendix. VSF also has the ability to generate “anti-aesthetics” (unconventional, including abstract) art. Details of these results are provided in Appendix Figure 18 and Figure 17.
6 Discussion
6.1 Trade Off Curve
To systematically evaluate how effectively each model balanced positive prompt adherence, negative prompt adherence, and image quality, we conducted a hyperparameter sweep across each model. Specifically, we performed 66 runs for VSF and 287 runs for NAG, and 10 runs for NASA, with respect to their hyperparameter counts (2 for VSF, 3 for NAG, and 1 for NASA). A random sweep was executed besides for NASA, on which a single variable “grid” search is used, and evaluations were conducted using LLaMA, following the same criteria as previously described. Due to the large volume of runs, we limited our evaluation to the first 100 prompts with a single generation seed, potentially resulting in minor differences from earlier outcomes. Results are shown in Figure 6 Left.
Figure 6: (Left) Trade-off plot of positive-negative score and quality-negative score. Both axies follows “higher is better.” (Right) Trade-off plot of the ablation study.
From both plots, we observe that as the negative score increases, NAG and NASA both exhibit a significantly steeper and earlier decline compared to VSF in both positive and quality scores. In terms of positive score, VSF maintains scores above 90 even when the negative score rises to approximately 60. Regarding image quality, VSF similarly retains scores above 90 until a negative score of around 60, after which quality declines. In contrast, NAG and NASA both experience a sharp and early decline, with their quality score rapidly dropping to nearly 60 even before the negative score reaches 50. Keep in mind that a quality score under 90 means the image is already degraded and if an image is rated 60, it is usually completely distorted. See Figure 11 in the Appendix for example.
Additionally, VSF demonstrates a broader operational range in negative scores. When necessary, it can achieve negative scores exceeding 70 while still preserving acceptable positive prompt adherence and image quality. Conversely, NAG and NASA become unacceptable in quality at negative scores below 50, limiting their practical effectiveness.
@@ 6.2 Attention Maps
Since our proposed method performs adaptive steering based on a negative attention map, we visualize the attention maps generated during the diffusion process in Figure 4. Extracting the full attention maps is difficult because efficient implementations, such as FlashAttention, do not explicitly store these maps, and storing and computing them will require a large amount of memory. Therefore, we computed only the unnormalized attention values between the image tokens and negative prompt tokens. Figure 4 demonstrates that when the scale is set to 0, umbrellas appear, whereas setting the scale to 3 effectively avoids them. As indicated in the attention maps, image tokens corresponding to regions where umbrellas might exist (e.g., above human heads) exhibit higher attention toward the negative prompt tokens. Specifically, in steps 4 and 5, regions above the individuals on the left and right show strong negative attention, aligning with areas visually identified as umbrellas in $\alpha=0$ image. In the final image, these highlighted regions no longer contain umbrellas, confirming that our method effectively suppresses the presence of undesired objects at specific locations.